The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu and Dominik J. Bomans

MCnews@astro.uiuc.edu

http://www.astro.uiuc.edu/mcnews/MCnews.html

Number: 5 June 20, 1996

Table of Contents

News and Views..............1 Paper Abstracts..............2 Job Advertisement..........6

News and Views

IAU Symposium on the Magellanic Clouds?

The 188th AAS meeting was held last week, and the Magellanic Clouds session went smoothly with a large turnout. We would like to thank all the speakers for their excellent talks and the participants for their interest and support.

During the meeting we discussed with a large number of participants about the idea of holding a major conference on the Magellanic Clouds in the near future. It is unanimously agreed that such a meeting is overdue, especially since vast amounts of new results have been produced by the Hubble Space Telescope, the ROSAT X-ray satellite, the Compact Array of Australia Telescope, and the advanced optical/IR detectors on ground-based telescopes.

We would like to solicit suggestions on an international meeting on the Magellanic Clouds. To start the discussion, a group of us MC researchers in the midwest (Illinois - Chu, Bomans, Dickel; Michigan - Smith; Minnesota - Dickey; Wisconsin - Wakker and Gallagher) propose an IAU symposium to be held in a nice resort area around the Great Lakes in the midwest of the US in 1998. Please send comments and suggestions to mcnews@astro.uiuc.edu. We look forward to your input. Thank you.

You-Hua Chu & Dominik Bomans

Paper Abstracts

Hubble Space Telescope Observations of Oxygen-rich Supernova Remnants in the Magellanic Clouds. I. Narrow-band Imaging of

N132D in the LMC^{1}

Jon A. Morse^{2,3}, William P. Blair^{4}, Michael A. Dopita^{5}, John P. Hughes^{6}, Robert P.

Kirshner^{6}, Knox S. Long^{2}, John C. Raymond^{6}, Ralph S. Sutherland^{5,7}, P. Frank Winkler^{8}

1 Based on observations with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555. 2 Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218 3 Center for Astrophysics and Space Astronomy, Department of Astrophysical, Planetary and Atmospheric Sciences, University of Colorado, Campus Box 389, Boulder, CO 80309 4 Department of Physics and Astronomy, Johns Hopkins University, Charles and 34th Streets, Baltimore, MD 21218 5 Mt. Stromlo and Siding Spring Observatories, Australian National University, Private Bag, Weston Creek Post Office, ACT 2611, Australia 6 Harvard-Smihsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138 7 JILA, University of Colorado and National Institute of Standards and Technology, Campus Box 440, Boulder, CO 80309 8 Department of Physics, Middlebury College, Middlebury, VT 05753

We present Hubble Space Telescope (HST) WFPC2 images of the young, oxygen-rich supernova remnant N132D in the Large Magellanic Cloud in the [O III]λ5007, [O II]λλ3727 and [S II]λλ6724 emission lines, plus a visible continuum band. The ∼ 0.1′′ resolution of HST allows us to analyze structure within N132D at scales comparable to ground-based images of nearby remnants such as Cas A. The oxygen-rich filaments (seen in [O III] and [O II] emission) are easily distinguished from shocked circumstellar clouds (seen in all three emission lines). Characteristic knots sizes and filament widths are ∼ 0.2′′ − 0.5′′ (∼ 1.5 − 4 × 10^{17} cm), at least partially resolved in our images. For the first time, we discern ionization structure within the shocked interstellar clouds and filaments, reminiscent of filaments in galactic SNRs that are thought to arise from recent shock/interstellar cloud encounters where the full cooling and recombination zone has not yet formed throughout the structure. Conversely, there appears to be fairly uniform ionization in the O-rich filaments. We also observe highly ionized, diffuse emission extending all the way around the outer edges of the remnant, as previously seen in ground-based studies. We compare our HST optical emission-line images of N132D to the ROSAT HRI soft X-ray image and find that the X-ray emission is closely traced optically by shocked circumstellar clouds; the O-rich filaments do not appear to be emitting X-rays. The limb-brightened X-ray shell aligns remarkably precisely along the inner edge of the diffuse outer rim. The diffuse emission probably represents ambient molecular gas that is being dissociated and photoionized by high-energy photons originating in the fast SNR shocks.

We have used the MAPPINGS II code to model the X-ray producing main blast wave as an ∼ 800 km s^{−1} shock moving into a circumstellar medium of density N_{0} ≈ 3 cm^{−3}. The preshock ISM density of N_{0} ≈ 3 cm^{−3} inferred from our shock model agrees with the preshock density estimated from the [O III]λ5007 surface brightness in the diffuse outer rim. We also model the [O III]/[S II] emission-line ratios in the photoionized outer rim as a function of distance ahead of the main blast wave using the radiation field produced by fast shocks in the remnant as the ionizing source. According to our models, the ionizing radiation from the ∼ 800 km s^{−1} main blast wave is not sufficient to generate the observed optical emission from the outer rim precursor. An extra source of EUV photons appears necessary to achieve the observed ionization of the preshock gas. We estimate that EUV photons produced in slower shocks moving into QSF-type clouds around the outer rim can probably contribute enough ionizing photons to achieve the observed ionization. Radiative shocks moving through the O-rich filaments may also be an important source of EUV photons. Accepted by The Astronomical Journal (August 1996 issue) For preprints, contact morsey@casa.colorado.edu Preprints are also available via the World Wide Web at http://casa.colorado.edu/ morsey/

Molecular abundances in the Magellanic Clouds: I. A multiline study of five cloud cores

Y.-N. Chin^{1,2}, C. Henkel^{3}, J.B. Whiteoak^{4}, T.J. Millar^{5}, M.R. Hunt^{6}, and C. Lemme^{3,7}

1 Radioastronomisches Institut der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2 Institute of Astronomy and Astrophysics, Academica Sinica, P.O.Box 1-87 Nankang, 115 Taipei, Taiwan 3 Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, D-53121 Bonn, Germany 4 Paul Wild Observatory, Australia Telescope National Facility, CSIRO, Locked Bag 194, Narrabri NSW 2390, Australia 5 Department of Physics, UMIST, P O Box 88, Manchester M60 1QD, United Kingdom 6 University of Western Sydney Nepean, P.O. Box 10, Kingswood, NSW 2747, Australia 7 Institut für Planetenerkundung, DLR, Rudower Chaussee 5, D-12489 Berlin, Germany

Nine H ii regions of the LMC were mapped in ^{13}CO(1–0) and three in ^{12}CO(1–0) to study the physical properties of the interstellar medium in the Magellanic Clouds. For N113 the molecular core is found to have a peak position which differs from that of the associated H ii region by 20′′. Toward this molecular core the ^{12}CO and ^{13}CO peak T_{MB} line temperatures of 7.3 K and 1.2 K are the highest so far found in the Magellanic Clouds. The molecular concentrations associated with N113, N44BC, N159HW, and N214DE in the LMC and LIRS 36 in the SMC were investigated in a variety of molecular species to study the chemical properties of the interstellar medium. I(HCO^{+})/I(HCN) and I(HCN)/I(HNC) intensity ratios as well as lower limits to the I(^{13}CO)/I(C^{18}O) ratio were derived for the rotational 1–0 transitions. Generally, HCO^{+} is stronger than HCN, and HCN is stronger than HNC. The high relative HCO^{+} intensities are consistent with a high ionization flux from supernovae remnants and young stars, possibly coupled with a large extent of the HCO^{+} emission region. The bulk of the HCN arises from relatively compact dense cloud cores. Warm or shocked gas enhances HCN relative to HNC. From chemical model calculations it is predicted that I(HCN)/I(HNC) close to one should be obtained with higher angular resolution (≤ 30′′) toward the cloud cores. Comparing virial masses with those obtained from the integrated CO intensity provides an H_{2} mass-to-CO luminosity conversion factor of 1.8 × 10^{20} mol cm^{−2} (K km s^{−1})^{−1} for N113 and 2.4 × 10^{20} mol cm^{−2} (K km s^{−1})^{−1} for N44BC. This is consistent with values derived for the Galactic disk. Accepted by Astronomy and Astrophysics For preprints, contact einmann@biaa21.biaa.sinica.edu.tw

Probing the ISM of the Superbubble LMC2 in the Large Magellanic Cloud: I. Ti II and Ca II absorption lines

Adeline Caulet^{1} and Ray Newell^{2}

1 ST-ECF, ESO, Karl–Schwarzschild–Str.2, Garching bei München, D–85748 Germany 2 Pomona College, Dept. of Physics, 610 N. College Ave., Claremont CA, 91711–6348

The optical interstellar absorption lines of Ti II and Ca II have been observed at high resolution towards seven OB supergiant stars in the Large Magellanic Cloud. The star lines of sight sample the disks and haloes of both the Galaxy and the LMC, H I clouds in the LMC, and the filamentary superstructure of ionized gas known as the Superbubble LMC2 near the 30 Doradus Nebula.

We have measured the velocities, equivalent widths, column densities, and Doppler–widths of Ca II and Ti II absorption lines detected over the whole range of velocities from −30 to +360 km s^{−1} between the Galaxy and the LMC. There is interstellar absorption at velocities measured for the Hα emission line gas of the superbubble and its surrounding giant H II regions, and for the H I gas components towards this region of the LMC. Absorbing gas is also found at intermediate velocities between 150 and 220 km s^{−1}, possibly identified with superbubble gas expanding into the LMC halo. The velocities above 150 km s^{−1} can be interpreted as three-dimensional motions of superbubble gas expanding in the surrounding medium in the framework of two existing representations of the large scale distribution of H I LMC gas. The first picture consists of overlapping H I sheets at separate velocities; the interpretation of velocity components follows our previous kinematical study of the Hα LMC2 filaments. The second picture consists of two main components, the rotating LMC disk (D) and a lower velocity H I component (L); the correlations between Hα, H I and absorption velocities suggest that LMC2 is expanding into the H I component L; perhaps, its origin comes from cloud collisions between the components L and D.

The remarkable extended absorption wings found between 150 and 220 km s^{−1} and the high velocity absorption components at ∼ 300 km s^{−1} are located in front of the H I components L and D; they prove the existence of upward and downward motions within the three-dimensional superbubble. Therefore LMC2 is an example where galactic fountains, superwinds and falling high velocity clouds participate to the complex dynamics of superbubbles. Accepted by The Astrophysical Journal (July 1) For preprints, contact acaulet@eso.org

SMC13 type supersoft sources

P. Kahabka^{1} and E. Ergma^{2}

1 Astronomical Institute and Centet for High Energy Astrophysics, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands 2 Physics Department, Tartu University, Ülikooli 18, EE2400 Tartu Estonia

It is proposed that the observational data for the supersoft X-ray source 1E0035.4-7230 in the SMC may be explained in the framework of the standard cataclysmic variable evolution. The fulfilling of several conditions are needed: (1) The white dwarf has a low mass (≈ 0.6 -0.7 M_{⊙}) and rather thick helium buffer layer on top of the C-O core. (2) After the common envelope phase and before the accretion stage the white dwarf has cooled down ( t_{c} ≈ 10^{8} − 10^{9} years). (3) After Roche lobe filling when the binary evolution is driven by magnetic braking, the mass accretion rate is ≈ 10^{−8} − 10^{−9} M_{⊙}/yr, the accumulation of the matter at the surface of the cold white dwarf leads to nova outbursts and all accreted matter will be ejected. (4) The accretion will gradually heat up the cool white dwarf and after reaching a steady state equilibrium higher temperature, flashes have become more mild and no longer all accreted matter is lost in the flash. At the current stage we are observing this system after a mild (due to the hot white dwarf condition) hydrogen flash in the phase of residual hydrogen burning. We would like to call these new type of objects SMC 13 type supersoft sources. Submitted to Astronomy and Astrophysics, research note For preprints, contact ptk@astro.uva.nl

Fundamental parameters of Wolf-Rayet stars VI. Large Magellanic Cloud WNL stars

P.A. Crowther and L.J. Smith

Department of Physics and Astronomy, University College London, Gower Street, London, WC1E 6BT, UK

We present a detailed, quantitative study of late WN (WNL) stars in the LMC, based on new optical spectroscopy (AAT, MSO) and the Hillier (1990) atmospheric model. In a previous paper (Crowther et al. 1995a), we showed that 4 out of the 10 known LMC Ofpe/WN9 stars should be reclassified WN9–10. We now present observations of the remaining stars (except the LBV R127), and show that they are also WNL (WN9–11) stars, with the exception of R99. Our total sample consists of 17 stars, and represents all but one of the single LMC WN6–11 population and allows a direct comparison with the stellar parameters and chemical abundances of Galactic WNL stars (Crowther et al. 1995b; Hamann et al. 1995a). Previously unpublished ultraviolet (HST-FOS, IUE-HIRES) spectroscopy are presented for a subset of our programme stars.

We find observational evidence for lower metallicities in LMC WNL stars compared to the Galaxy, though this is not reflected in their stellar properties. For Galactic and LMC stars we find: (i) a similar range in temperature and luminosity, in contrast to evolutionary predictions; (ii) comparable wind performance values (Ṁ v_{∞} /[L/c]) and hydrogen composition, with a broad correlation between increasing helium content and wind performance number; (iii) a general trend to lower wind velocities at lower stellar temperature, with possibly slower winds for LMC WN9–11 stars. Some 30 Dor WNL stars show exceptional properties: Brey 89 (HD 38282, WN6h) has the highest luminosity (log (L/L_{⊙})∼6.25) and mass-loss rate (log (Ṁ/(M_{⊙} yr^{−1})) ∼−3.6) known for any WR star, while Brey 80 (R135, WN7h) has an enormous wind performance number of 50. The observed physical properties of our sample of LMC WNL stars supports the Crowther et al. (1995c) evolutionary scheme for Galactic stars, in that the most massive O stars, exclusive to 30 Dor, evolve directly to O3 If/WN6 and subsequently WN6–7 stars (e.g. Brey 89), without passing through an intermediate LBV phase. In contrast, lower initial mass stars evolve through a LBV phase, encompassing a WN9–11 stage (e.g. BE294), with WN8 stars being their immediate successors. Accepted by Astronomy & Astrophysics Preprints available via anonymous ftp - please contact pac@star.ucl.ac.uk

Be star surveys with CCD photometry. II. NGC 1818 and its neighbouring cluster in the LMC

E.K. Grebel

Accepted by Astronomy & Astrophysics, Main Journal (June 7, 1996) abstract in MCnews #3 For preprints, contact grebel@astro.uni-wuerzburg.de The paper is available at http://www.astro.uiuc.edu/∼grebel/publications.html

Job Advertisement

Postdoctoral Research Position

Columbia University Columbia Astrophysics Laboratory 538 W. 120th St. New York, NY 10027

Attn: Arlin Crotts Phone: (212) 854-7899 FAX: (212) 854-8121 email: arlin@eureka.phys.columbia.edu

We are searching for a postdoctoral scientist to aid in the research on the circumstellar and interstellar environment of Supernova 1987A. The work would include completed, ongoing and planned observations of SN 1987A and its surroundings including data from IUE, UIT, ISO, HST and various ground-based observatories, plus preparation for observations with future ground, airborne and spacebased instruments. Topics of concern include mapping the circumstellar nebula of the SN and the interstellar medium of the 30 Doradus region using light echoes, detailing the composition, density and velocity structure of the circumstellar envelope, analyzing the polarimetric and photometric properties of surrounding dust, measuring the reflected ultraviolet flux from SN shock breakout, and observing the interaction between the SN ejecta and circumstellar envelope and the formation of Supernova Remnant 1987A. We are looking for excellent candidates practiced in some of the above observational techniques and knowledgeable concerning circumstellar and interstellar processes. The person selected will be encouraged to study the above problems as well as to develop his or her own related investigations, with support for observation, publication and travel to professional meetings. The position is available for one year with the possibility of extension to up to three years, with a start date in early 1997 subject to the initiation of funding. Applicants should provide a curriculum vitae, summary of research, and names of three colleagues willing to write letters of reference, by 1 July 1996 to the address above. Columbia University is an Affirmative Action/Equal Opportunity Employer.